# Task-Agnostic Continual Prompt Tuning with Gradient-Based Selection and Decoding

### v1

@@ Abstract
Prompt-based continual learning (CL) offers a parameter-efficient way to adapt large language models (LLMs) across task sequences. However, most existing methods assume task-aware inference and maintain a growing list of task-specific prompts, which limits scalability and hides latent forgetting. In this work, we introduce GRID, a unified framework that addresses two key limitations: (1) latent forgetting under task-agnostic inference, and (2) prompt memory explosion as task sequences grow. GRID integrates a task-aware decoding mechanism that improves backward transfer by leveraging representative inputs, automatic task identification, and constrained decoding. Additionally, we propose a gradient-based prompt selection strategy that compresses less informative prompts into a single aggregated representation, enabling scalable and memory-efficient lifelong learning. Extensive experiments across short-sequence, long-sequence, and negative transfer benchmarks show that GRID significantly improves backward transfer, achieves competitive forward transfer, and reduces forgotten tasks by up to 80%, outperforming state-of-the-art methods on T5 and Flan-T5 backbones. Source code is available here.
@@ 1 Introduction
Continual learning (CL) [37] enables models to learn from a sequence of tasks without retraining from scratch. CL systems build on prior knowledge while adapting to new tasks, which is crucial in dynamic real-world settings. Recent advancements, especially in NLP [42, 33], have focused on three paradigms: regularization-based methods [20, 17], rehearsal-based methods [28, 36], and architecture-based methods [38, 10]. While rehearsal-based methods are effective, they are impractical in privacy-sensitive scenarios. With the increasing complexity of pretrained models [41], full model finetuning has become infeasible. This has led to the rise of parameter-efficient finetuning (PEFT) techniques [9], such as prompt tuning (PT) [18], which adapts large models by training only a small set of soft prompts, requiring less than 0.01% of the model’s parameters.
@@
Building on this, continual prompt tuning (CPT) extends PT to the CL setting by enabling models to learn task-specific prompts sequentially without modifying the base model [45]. Recent advancements in prompt tuning-based continual learning (CPT) address key challenges like catastrophic forgetting and forward knowledge transfer. [44] introduced a dual prompt framework with a shared and task-specific prompt, though this approach faces limitations in retaining knowledge from past tasks. To mitigate these issues, ProgPrompt [27] maintains a list of prompts for each task, progressively adding new ones while retaining old ones. Coda-prompt [35] extended this by proposing prompt set expansion using weighted summation to ensure efficient knowledge transfer. A substantial body of work has focused on improving CPT’s ability to retain and transfer knowledge, using techniques such as prompt concatenation [27], parameter sharing [45], and weighted prompt summation [15].
Although significant attention has been given, there are still two key challenges in prompt-based CL: (1) latent forgetting under task-agnostic inference and (2) scalability issues from growing prompt queues. First, recent work in prompt-based continuous learning often assumes that assigning a separate prompt to each task eliminates catastrophic forgetting [22], since the model’s core parameters and previously learned prompts remain untouched [18]. Following this assumption, approaches such as Progressive Prompts [27], Q-Tuning [13], and SHLPT [47] report zero Backward Transfer (BWT), suggesting that prior task knowledge is fully preserved. However, these claims are based on task-aware inference, where the identity of the task is known and used to select the corresponding prompt. In real-world scenarios, this assumption often fails: not only is task identity typically unavailable at inference time, but storing a growing number of task-specific prompts becomes increasingly impractical in continual learning settings with an increasing number of tasks. Under such task-agnostic setups, the model must generalize across all previously learned tasks without relying on task-specific prompts. In this setting, our extensive experiments show that the performance on earlier tasks degrades after training on new ones, with the model frequently generating incorrect or ambiguous labels. This issue is particularly problematic for encoder-decoder models like T5, which generate output labels as text. Without clear task guidance, the model may produce label words from unrelated tasks it has seen during pretraining or prior learning. These observations point to a form of latent forgetting that is obscured by task-aware evaluation. In particular, this limitation is also acknowledged in recent works like [13], which find that despite having strong forward transfer, it is not capable of supporting backward knowledge transfer or task-agnostic inference.
The second challenge in prompt-based continual learning lies in addressing the scalability issues in time and memory as the prompt queue grows. Approaches such as Progressive Prompts [27] train and retain a separate soft prompt for each task, resulting in a prompt list that grows linearly with the number of tasks. As task sequences grow, this leads to $\mathcal{O}(N)$ prompt memory and $\mathcal{O}(N^{2})$ inference complexity in transformer-based models. The recent Q-Tuning method [13] addresses this challenge by maintaining a bounded-size prompt queue and discarding less informative prompts using a principal component analysis (PCA)-based eviction strategy. However, it still encounters a trade-off between performance and capacity, and incurs non-negligible computational overhead due to the singular value decomposition (SVD) required for each PCA-based eviction.
In this paper, we aim to address these challenges with the following key contributions:
  • We are the first to reveal that existing prompt-based CL methods reporting $\mathbf{0}$ BWT rely on task-aware inference, which masks latent forgetting in task-agnostic settings. We also identify scalability limitations due to prompt memory growing linearly with task count.
  • We propose GRID (Gradient-based prompt selection with Representative samples selection, task Identification, and Decoding), a unified framework with two core components.
@@
    • A task-aware decoding mechanism that leverages representative inputs and constrained decoding to ensure label consistency and improve backward knowledge retention (BWT) without needing task IDs.
@@
    • A gradient-based prompt selection mechanism that dynamically evaluates the usefulness of prompts and merges less informative ones, reducing storage by 66.7% while maintaining performance.
@@
  • We conduct extensive experiments on encoder-decoder models (T5 and Flan-T5) across short, long, and negative transfer sequences. GRID achieves up to 54% BWT improvement, reduces number of forgotten tasks by 80%, and matches or outperforms state-of-the-art baselines such as ProgPrompt, and SHLPT, under realistic task-agnostic settings.
2 Related Work
Continual Learning. Continual learning (CL) involves learning from a sequence of tasks without full access to previous task data, aiming to preserve prior knowledge and enable positive transfer. A key challenge is catastrophic forgetting [22], where updates to model parameters on new data erode earlier knowledge. CL strategies are categorized into three approaches: memory-based methods store and replay past task data to mitigate forgetting [34, 3]; regularization-based methods penalize deviations from important parameters to retain knowledge without stored data [17, 51]; and architecture-based methods expand the model by adding new components for each task [32, 50]. These methods face scalability issues in large pretrained models, motivating the development of parameter-efficient CL techniques using lightweight components like prompts and adapters [48, 31].
Continual Prompt Tuning. Prompt tuning [18, 19, 12, 43, 14] is a parameter-efficient method for adapting large language models (LLMs) to new tasks by learning a small set of continuous vectors, known as soft prompts, which are prepended to the input tokens. Unlike traditional finetuning, which modifies all model parameters, prompt tuning optimizes only the prompt parameters, leaving the rest of the model unchanged. This approach has been shown to match or surpass the performance of full model finetuning, offering substantial savings in computational cost and memory usage. Building on this, continual prompt tuning [53, 49, 11, 44, 27] extends prompt tuning to the continual learning (CL) setting, where models must adapt to evolving task sequences [18, 12]. However, existing continual prompt tuning methods face challenges, such as the reliance on large memory buffers to prevent forgetting [53, 11], which are impractical in privacy-sensitive or resource-constrained settings. Additionally, task-specific prompts still suffer from latent forgetting in task-agnostic inference scenarios [13], and strategies that incrementally grow prompt representations [27, 44, 35] struggle with scalability. As the number of tasks increases, the size of the prompt grows, leading to efficiency bottlenecks in both training and inference.
Gradient-Based Data and Parameter Selection. Gradient-based strategies have been extensively studied in data selection and coreset construction [2], where the goal is to identify a representative or influential subset of training data [4, 30]. Methods such as Coreset Selection [16, 23] and Gradient Matching [2] leverage the similarity or norm of training gradients to preserve performance while reducing dataset size. Inspired by these ideas, recent works have begun to explore gradient-based criteria for parameter-efficient tuning. In the context of prompt learning, Q-Tuning [13] proposes managing a fixed-size prompt queue but does not use explicit gradient scoring to retain or merge prompts. Our work draws inspiration from the coreset literature and extends it to the prompt space by dynamically evaluating gradient norms of prompt embeddings, allowing us to merge less informative prompts and improve scalability under long task sequences.
3 Background and Motivation
3.1 Prompt-based Continual Learning and Task-Agnostic Inference
@@
Problem Setup. We consider a continual learning (CL) setting in which a model encounters a sequence of $N$ tasks $\mathcal{T}=\{T_{1},T_{2},\dots,T_{N}\}$, where each task $T_{i}$ is associated with a labeled dataset $D_{i}=\{(x_{j},y_{j})\}_{j=1}^{|D_{i}|}$. Here, $x_{j}$ denotes an input instance and $y_{j}\in\mathcal{Y}_{i}$ is the corresponding label. The model is built upon a pretrained encoder-decoder language model $f(\cdot;\theta)$, whose parameters $\theta$ are kept fixed throughout learning. Rather than finetuning $\theta$, we adapt the model to each task $T_{i}$ by learning a soft prompt $\mathbf{p}_{i}\in\mathbb{R}^{l\times d}$, where $l$ denotes the prompt length and $d$ the embedding dimension. After observing tasks $\{T_{1},\dots,T_{t-1}\}$, we maintain a pool of learned prompts $\mathcal{P}=\{\mathbf{p}_{1},\dots,\mathbf{p}_{t-1}\}$, which serves as a memory of past task adaptations.
@@
Inference Settings. We consider two commonly studied inference modes in continual learning:
1) Task-Aware Inference: The task identity is known at inference time. The model selects the corresponding prompt $\mathbf{p}_{i}$ from the prompt pool $\mathcal{P}$ to perform prediction on task $T_{i}$.
2) Task-Agnostic Inference: The task identity is unknown at inference time. As a result, the model cannot rely on task-specific prompt selection. Instead, it performs prediction using the entire available prompt pool $\mathcal{P}=\{\mathbf{p}_{1},\dots,\mathbf{p}_{t-1}\}$, which may optionally include an aggregated prompt derived from previous filtering stages. The concatenated prompt sequence is prepended to the input and passed to the model. Formally, the prediction is defined as: $\hat{y}=f(x;\mathcal{P})$.
The task-agnostic evaluation objective then becomes: $\mathcal{L}_{\text{TA}}=\sum_{T_{k}\in\mathcal{T}_{\text{past}}}\mathbb{E}_{(x,y)\sim D_{k}}\left[\ell(f(x;\mathcal{P}),y)\right].$ This formulation evaluates how well the concatenated prompt pool enables generalization to earlier tasks without needing to retrieve or know their corresponding individual prompts. This setting appears in two important continual learning scenarios: 1) Online learning, where a new task arrives and the model must evaluate or detect alignment with previously seen tasks [1]. 2) Retroactive evaluation, where a model trained on a task sequence is later evaluated on older tasks without access to their individual prompts [5, 8].
3.2 Discussion of Challenges in Existing Prompt-based CL
Despite recent progress, prompt-based continual learning methods suffer from critical limitations: Label Drift and Hallucination. In task-agnostic settings, encoder-decoder models generate free-form text outputs without explicit constraints, often resulting in label drift—predictions biased toward labels from unrelated tasks or the pretraining corpus. While the encoder may capture relevant contextual features, the decoder can produce hallucinated outputs that are semantically invalid or inconsistent with the intended task. We provide qualitative examples of such failures in Table 1.
Table 1: Qualitative comparison under task-agnostic inference. Blue indicates label drift, where Progressive Prompts predict semantically related but incorrect labels. Red highlights hallucination, where the model generates invalid outputs not part of the task’s label set.
@@ 3.2 Discussion of Challenges in Existing Prompt-based CL
Inconsistent Label Mappings. Prior methods such as Progressive Prompts [27] and SHLPT [47] rely on human-specified mappings (e.g., 0 $\rightarrow$ "negative") to convert numeric labels into text. These mappings are often inconsistent across datasets and introduce semantic misalignment that hinders generalization across tasks. Our Appendix (e.g., Table 21) further highlights the cases where such mappings cause contradictory model behavior.
Lack of Task Awareness. Current methods do not attempt to infer or utilize the underlying task structure. As a result, tasks with similar semantics (e.g., different sentiment classification datasets) are treated independently, missing opportunities for generalization or shared label space alignment. This leads to redundant prompt representations and missed transfer potential.
Unbounded Prompt Growth. Approaches like Progressive Prompts [27] maintain a separate prompt for each task and concatenate them during inference. This causes prompt pool size to grow linearly with the number of tasks, increasing both memory usage and inference time. Moreover, prompt concatenation can unpredictably affect decoding behavior, especially under task-agnostic evaluation.
These theoretical and practical limitations highlight the need for a more structured and scalable framework for prompt-based continual learning under task-agnostic conditions. We provide empirical evidence in Table 1 showing that label hallucination, and semantic drift persists even when the encoder produces reasonable representations. These findings motivate our task-aware input construction and prompt queue management strategies. Furthermore, several open questions surrounding the inference-time behavior of progressive prompts remain unresolved [^1] See discussion at this issue on GitHub., which our work seeks to address.
@@ 4 Method
Reformulation. To address the dual challenges of unstable decoding under task-agnostic inference and unbounded prompt pool growth, we propose a unified framework called GRIDGradient-based prompt selection with Representative sample selection, task Identification, and constrained Decoding. GRID integrates two complementary components: 1) A task-aware input pipeline that enhances backward retention and output consistency by selecting representative samples, performing task identification, and applying constrained decoding; 2) A gradient-based prompt scoring mechanism that reduces prompt pool size by identifying and merging less informative prompts while preserving relevant task knowledge. Together, these components enable scalable and reliable continual learning under task-agnostic inference. See Figure 1 for the pipeline and Appendix A for full algorithms.
4.1 Task-Aware Input Construction for Stable Decoding
@@
1) Representative Input Sampling. To construct a compact yet informative training subset, we select $k=1000$ representative samples per class via clustering, improving upon the random sampling strategy used in prior work [27]. The data set $\mathcal{D}=\{(x_{j},y_{j})\}$ is partitioned by label $y\in\mathcal{Y}$, and the sentences $\mathbf{e}_{j}=f_{\text{embed}}(x_{j})$ are calculated using a pre-training embedding model (all-MiniLM-L6-v2; 29). K-Means clustering is performed within each class to ensure intra-class diversity, and the top $k/C$ samples closest to each cluster center are selected based on cosine similarity: $\text{sim}(\mathbf{e},\mathbf{c})=\frac{\mathbf{e}\cdot\mathbf{c}}{\|\mathbf{e}\|\|\mathbf{c}\|}.$ This yields a balanced subset $\mathcal{D}_{\text{rep}}$ that spans the semantic space of each class. We empirically find that increasing beyond 1k samples per class, the accuracy either saturates or improves negligibly, indicating that larger sample sizes may not result in substantial gains in performance (See Appendix B Table 10), while clustering improves both data efficiency and generalization.
2) Task Identification. In many datasets, task categories are not explicitly provided, and labels often appear in non-descriptive formats (e.g., {0,1} or {choice1, choice2}) that are ambiguous without task context. To resolve this, we implement a hierarchical task identification module that infers the task type $t^{*}$ given a candidate label set $\mathcal{Y}_{i}$ and a sample input $x$. The process begins with rule-based heuristics, which match label tokens and input structure to predefined task templates (e.g., “positive/negative” for sentiment analysis, or premise-hypothesis pairs with “entailment/contradiction” for NLI). If this fails, we fallback to zero-shot task classification using a lightweight generative language model (e.g., Phi-3.5). The model is prompted to infer the task type and remap non-descriptive labels into meaningful textual tokens: $(t^{*},\tilde{\mathcal{Y}}_{i})=\textbf{LLM}(x,\mathcal{Y}_{i}).$ This remapping enables consistent semantic interpretation across tasks, making it possible to apply decoding constraints in a task-agnostic setting.
@@
4.2 Prompt Pool Compression via Gradient-Based Selection
Prompt-based continual learning methods typically store one prompt per task, leading to prompt pool growth of $\mathcal{O}(N)$. To enable scalable lifelong learning, we dynamically select and merge prompts based on their gradient relevance to the current task. Let $\mathcal{P}=\{\mathbf{p}_{1},\dots,\mathbf{p}_{t-1}\}$ denote the prompt pool before learning task $T_{t}$. For each prompt $\mathbf{p}_{j}\in\mathcal{P}$, we compute the average gradient norm of the loss $L$ over the new task data $D_{t}$:
$g_{j}=\frac{1}{|D_{t}|}\sum_{(x,y)\in D_{t}}\|\nabla_{\mathbf{p}_{j}}L(f(x;\mathbf{p}_{j}),y)\|_{2}.$ (1)
A high gradient norm indicates that the current task significantly alters the prompt, suggesting it contains distinct knowledge not yet captured—making it valuable for maintaining diversity. In contrast, prompts with low gradient norms align closely with the current task, indicating redundancy; hence, they are compressed to retain the potential for forward transfer while reducing memory footprint. Prompts are partitioned using a threshold:

### v3

# GRID: Scalable Task-Agnostic Prompt-Based Continual Learning for Language Models

@@ Abstract
Prompt-based continual learning (CL) provides a parameter-efficient approach for adapting large language models (LLMs) across task sequences. However, most existing methods rely on task-aware inference and maintain a growing set of task-specific prompts, which introduces two major challenges: (1) severe performance degradation on earlier tasks under task-agnostic inference, and (2) limited scalability due to prompt memory accumulation as task sequences grow. In this paper, we present GRID, a unified framework designed to address these challenges. GRID incorporates a decoding mechanism that enhances backward transfer by leveraging representative inputs, automatic task identification, and constrained decoding. Furthermore, it employs a gradient-guided prompt selection strategy to compress less informative prompts into a single aggregated representation, ensuring scalable and memory-efficient continual learning. Extensive experiments on long-sequence and negative transfer benchmarks show that GRID improves average accuracy and backward transfer, achieves competitive forward transfer, and substantially reduces prompt memory usage. Source code is available here.
@@ 1 Introduction
Continual learning (CL) (Van de Ven & Tolias, 2019) enables models to learn from a sequence of tasks without retraining from scratch. CL systems build on prior knowledge while adapting to new tasks, which is crucial in dynamic real-world settings. Recent advancements, especially in NLP (Wang et al., 2024; Satapara & Srijith, 2024), have focused on three paradigms: regularization-based methods (Li & Hoiem, 2017; Kirkpatrick et al., 2017), rehearsal-based methods (Rebuffi et al., 2017; Sun et al., 2019), and architecture-based methods (Veniat et al., 2020; Douillard et al., 2022). While rehearsal-based methods are effective, they are impractical in privacy-sensitive scenarios. With the increasing complexity of pretrained models (Wang et al., 2023a), full model finetuning has become infeasible. This has led to the rise of parameter-efficient finetuning (PEFT) techniques (Ding et al., 2023), such as prompt tuning (PT) (Lester et al., 2021), which adapts large models by training only a small set of soft prompts, requiring less than 0.01% of the model’s parameters.
@@
Building on this, continual prompt tuning (CPT) extends PT to the CL setting by enabling models to learn task-specific prompts sequentially without modifying the base model (Wang et al., 2022b). Recent advancements in prompt tuning-based continual learning address key challenges like catastrophic forgetting and forward knowledge transfer. Wang et al. (2022a) introduced a dual prompt framework with a shared and task-specific prompt, though this approach faces limitations in retaining knowledge from past tasks. To mitigate these issues, ProgPrompt (Razdaibiedina et al., 2023) maintains a list of prompts for each task, progressively adding new ones while retaining old ones. SHLPT (Wu et al., 2024) extended this by proposing an attention-based similarity estimator to compose similar prompts for initialization while contrastively regularizing against dissimilar tasks, mitigating negative transfer.
Although prompt-based CL has received significant attention, two main challenges remain:
  • $\bullet$
    First, while major prompt-based CL methods such as Progressive Prompts (Razdaibiedina et al., 2023), Q-Tuning (Guo et al., 2024), and SHLPT (Wu et al., 2024) report zero forgetting, thereby fully preserving prior knowledge, they depend on task-aware inference, where task identities are explicitly provided to retrieve the correct prompt for each previous task. In practice, however, task IDs are often unavailable during inference (Wang et al., 2024; Liang & Li, 2023).
  • $\bullet$
    Second, most prompt-based CL methods struggle to scale efficiently in both time and memory as the prompt queue expands. For example, Progressive Prompts (Razdaibiedina et al., 2023) and SHLPT (Wu et al., 2024) assign a dedicated soft prompt to each task, leading to memory usage that grows linearly with the number of tasks. More recent approaches attempt to mitigate this issue by using PCA-based eviction (Guo et al., 2024) or by continually updating the same set of prompts over time (Wang et al., 2022b). However, these solutions either incur significant computational overhead (e.g., repeated SVD during eviction) or suffer from redundant prompt accumulation due to the lack of effective pruning or merging mechanisms.
The above challenges motivate us to investigate an important but underexplored task-agnostic scenario:
  • The task identity is unknown at inference time, and the prompt pool has a fixed capacity. In this setting, the model cannot (1) guarantee zero forgetting by relying on task-specific prompts, nor (2) indefinitely store all prompts as the task sequence grows. Consequently, the model must perform prediction using the entire available prompt pool, while prompt pruning or merging becomes necessary to maintain the prediction quality as well as the bounded size of the pool.
This paper focuses on the above setting and makes the following specific contributions.
  • $\bullet$
    We observe that existing prompt-based continual learning approaches often struggle in task-agnostic scenarios. Our experiments reveal that when task identities are unavailable at inference time, performance on earlier tasks degrades substantially after training on new ones, with models frequently producing incorrect or ambiguous outputs. This issue is especially pronounced for encoder–decoder architectures, which generate labels as free text: without explicit task cues, the model may output label words from unrelated tasks encountered during pretraining or earlier learning. We refer to this phenomenon as latent forgetting, denoting the degradation in performance on earlier tasks under task-agnostic evaluation. Similar observations have been reported in prior work (e.g., (Guo et al., 2024)), showing that forward transfer often persists while backward transfer suffers severely in the task-agnostic setting.
  • $\bullet$
    We introduce GRID, a unified framework that integrates constrained decoding with a principled gradient-guided prompt selection to address the above limitations:
@@
    • A decoding mechanism that leverages representative inputs and constrained decoding, an approach applied to control the output space of pre-trained LLMs, to ensure label consistency and improve backward transfer (BWT) without relying on task IDs.
@@
    • A gradient-guided prompt selection strategy that dynamically evaluates prompt usefulness and merges less informative prompts, significantly reducing memory usage while maintaining both forward and backward transfer performance.
  • $\bullet$
@@
    We conduct extensive experiments across long-sequence and negative transfer benchmarks. GRID improves BWT by up to 54%, reduces the number of forgotten tasks by 80%, and consistently matches or outperforms average accuracy of the state-of-the-art prompt-based CL baselines such as ProgPrompt and SHLPT under the task-agnostic conditions.
2 Background and Challenges
2.1 Problem Setup
@@
We consider a continual learning (CL) setting in which a model encounters a sequence of $N$ tasks $\mathcal{T}=\{T_{1},T_{2},\dots,T_{N}\}$, where each task $T_{i}$ is associated with a labeled dataset $D_{i}=\{(x_{j},y_{j})\}_{j=1}^{|D_{i}|}$. Here, $x_{j}$ denotes an input instance and $y_{j}\in\mathcal{Y}_{i}$ is the corresponding label. The model is built upon a pretrained encoder-decoder language model $f(\cdot;\theta)$, whose parameters $\theta$ are kept fixed throughout learning. Rather than finetuning $\theta$, we adapt the model to each task $T_{i}$ by learning a soft prompt $\mathbf{p}_{i}\in\mathbb{R}^{l\times d}$, where $l$ denotes the prompt length and $d$ the embedding dimension. After observing tasks $\{T_{1},\dots,T_{t-1}\}$, we maintain a pool of learned prompts $\mathcal{P}=\{\mathbf{p}_{1},\dots,\mathbf{p}_{t-1}\}$, which serves as a memory of past task adaptations.
@@
2.2 Task-Agnostic Inference Setting
Building on the aforementioned motivation, this paper focuses on the following task-agnostic inference setting:
Definition 1 (Task-Agnostic Inference).
The task identity is unknown at inference time. As a result, the model cannot rely on task-specific prompt selection. Instead, it performs prediction using the entire available prompt pool $\mathcal{P}=\{\mathbf{p}_{1},\dots,\mathbf{p}_{t-1}\}$, which may optionally include an aggregated prompt derived from previous filtering stages. The concatenated prompt sequence is prepended to the input and passed to the model. Formally, the prediction is defined as: $\hat{y}=f(x;\mathcal{P})$.
The task-agnostic evaluation objective becomes: $\mathcal{L}_{\text{TA}}=\sum_{T_{k}\in\mathcal{T}_{\text{past}}}\mathbb{E}_{(x,y)\sim D_{k}}\left[\ell(f(x;\mathcal{P}),y)\right].$ This formulation evaluates how well the concatenated prompt pool enables generalization to earlier tasks without needing to retrieve or know their corresponding individual prompts. This setting appears in two important CL scenarios: 1) Online learning, where a new task arrives and the model must evaluate or detect alignment with previously seen tasks (Aljundi et al., 2019a). 2) Retroactive evaluation, where a model trained on a task sequence is later evaluated on older tasks without access to their individual prompts (Chaudhry et al., 2018; de Masson D’Autume et al., 2019).
2.3 Challenges in Existing Prompt-based CL in Task-agnostic Setting
In this section, we discuss several challenges of existing methods in the task-agnostic case.
Label Drift and Hallucination. In task-agnostic settings, encoder–decoder models such as T5 generate free-form text outputs without explicit constraints, which often leads to two types of errors. The first is label drift, where the model produces semantically related but incorrect labels or syntactically invalid variants. The second is hallucination, where the decoder generates unseen or spurious labels due to the likelihood mass fragmenting over an expanding vocabulary. Qualitative examples of these failure cases are provided in Table 1.
Table 1: Qualitative comparison under task-agnostic inference. Blue indicates label drift, where Progressive Prompts predict syntactically incorrect labels. Red highlights hallucination, where the model generates invalid outputs not part of the task’s label set.
@@ 2.3 Challenges in Existing Prompt-based CL in Task-agnostic Setting
Inconsistent Label Mappings. Prior approaches such as Progressive Prompts (Razdaibiedina et al., 2023) and SHLPT (Wu et al., 2024) typically rely on human-specified mappings (e.g., 0 $\rightarrow$ “negative”) to convert numeric labels into text. Since these mappings can vary across datasets, they may introduce semantic inconsistencies that make cross-task generalization more challenging. Illustrative examples are provided in the Appendix (Table 19), showing how such differences can affect model behavior.
Lack of Task Awareness. Current methods often overlook the underlying task structure. As a result, even semantically related tasks (e.g., sentiment classification across different datasets) may be treated separately, which can limit opportunities for generalization and shared label alignment. This may lead to overlapping prompt representations and reduced transfer.
Prompt Growth. Under task-agnostic evaluation, the model makes predictions based on the entire prompt pool. Without mechanisms for pruning or merging, the prompt memory can gradually increase with possibly redundant prompt accumulation, which is not explicitly addressed in some existing prompt-based CL approaches such as Progressive Prompts (Razdaibiedina et al., 2023).
These challenges highlight the need for a more structured and scalable framework for prompt-based continual learning under task-agnostic conditions.
@@ 3 Method
To address the aforementioned challenges, we propose a unified framework called GRID: Gradient-based prompt selection with Representative sample selection, task Identification, and constrained Decoding. GRID integrates two complementary components: 1) An input pipeline that enhances backward retention and output consistency by selecting representative samples, performing task identification, and applying constrained decoding; 2) A gradient-based prompt scoring mechanism that reduces prompt pool size by identifying and merging less informative prompts while preserving relevant task knowledge. Figure 1 for the pipeline and Appendix A for full algorithms.
3.1 Input Construction for Stable Decoding
@@
1) Representative Input Sampling. To construct a compact yet informative training subset, we select $k$ representative samples per class via clustering, improving upon the random sampling strategy used in prior work (Razdaibiedina et al., 2023). The data set $\mathcal{D}=\{(x_{j},y_{j})\}$ is partitioned by label $y\in\mathcal{Y}$, and the sentences $\mathbf{e}_{j}=f_{\text{embed}}(x_{j})$ are calculated using a pre-trained embedding model (all-MiniLM-L6-v2; (Reimers & Gurevych, 2019)). K-Means clustering is performed within each class to ensure intra-class diversity, and the top $k/C$ samples closest to each cluster center are selected based on cosine similarity: $\text{sim}(\mathbf{e},\mathbf{c})=\frac{\mathbf{e}\cdot\mathbf{c}}{\|\mathbf{e}\|\|\mathbf{c}\|}.$ This yields a balanced subset $\mathcal{D}_{\text{rep}}$ that spans the semantic space of each class.
2) Task Identification. In many datasets, task categories are not explicitly provided, and labels often appear in non-descriptive formats (e.g., {0,1} or {choice1, choice2}) that are ambiguous without task context. To resolve this, we implement a hierarchical task identification module that infers the task type $t^{*}$ given a candidate label set $\mathcal{Y}_{i}$ and a sample input $x$. The process begins with rule-based heuristics, which match label tokens and input structure to predefined task templates (e.g., “positive/negative” for sentiment analysis, or premise-hypothesis pairs with “entailment/contradiction” for NLI). If this fails, we fallback to zero-shot task classification using a lightweight generative language model (e.g., Phi-3.5). The model is prompted to infer the task type and remap non-descriptive labels into meaningful textual tokens: $(t^{*},\tilde{\mathcal{Y}}_{i})=\textbf{LLM}(x,\mathcal{Y}_{i}).$ This remapping enables consistent semantic interpretation across tasks, making it possible to apply decoding constraints in the task-agnostic setting.
@@
3.2 Prompt Pool Compression via Gradient-Guided Selection
Prompt-based CL methods typically allocate one prompt per task, causing the prompt pool to grow as $\mathcal{O}(N)$. To achieve scalable long-horizon continual learning, we propose a dynamic selection–compression strategy based on gradient relevance. Let $\mathcal{P}=\{\mathbf{p}_{1},\dots,\mathbf{p}_{t-1}\}$ be the prompt pool before task $T_{t}$. For each $\mathbf{p}_{j}\in\mathcal{P}$, we compute the average gradient norm over new task data $D_{t}$:
$g_{j}=\tfrac{1}{|D_{t}|}\sum_{(x,y)\in D_{t}}\|\nabla_{\mathbf{p}_{j}}L_{t}(f(x;\mathbf{p}_{j}),y)\|_{2}.$ (1)
A large $g_{j}$ indicates that the task substantially updates $\mathbf{p}_{j}$, suggesting distinct knowledge worth preserving. Conversely, small $g_{j}$ values imply redundancy with the current task. Prompts are partitioned using the threshold